# Presenting a STEM Ways of Thinking Framework for Engineering Design-based Physics Problems

### v3

@@ 3.4 The Five WoT4EDP Elements
DBT Economic aspects, safety aspects, criteria and constraints, identify stakeholders, contraption dimensions, contraption mechanism, design limitations, feasibility study. [20, 21] Design thinking helps students develop problem-solving skills, creativity, empathy, collaboration, iterative thinking, and multiple perspectives. [53] Incorporate real-world, open-ended problems to encourage creative and user-centered solutions. Use iterative prototyping and reflective feedback sessions to help students refine ideas and increase empathy. Collaborate across disciplines to stimulate innovation. A focus on failure as a learning opportunity strengthens resilience. The ‘design thinking competency model’ [54, p.342] may be used as a guide.
@@
Based on our readings [69, p.4], we gather that the term ‘design thinking’ likely originated with Rowe, who used it as the title of his 1987 book [70], in which he aimed to provide a structured account of problem-solving processes in design. Since then, however, the meaning of the term design thinking has undergone several transformations and has drawn varied interpretations.
Haasi and Laakso [69, p.5] describe design thinking as comprising three key elements: (i) practices, including human-centered approaches, thinking by doing, visualizing, integrating multiple approaches, and collaboration; (ii) cognitive approaches, such as abductive reasoning, reflective reasoning, a holistic perspective, and integrative thinking; and (iii) a mindset characterized by exploration, tolerance for ambiguity, optimism, and a future-oriented outlook.
@@
While the debate over the usage of design thinking is thought-provoking, we find merit in the use of the term in our contexts. For operational purposes, one may view design thinking as a problem-solving approach that emphasizes creativity, user-centered solutions, and iterative processes. It involves understanding users’ needs, brainstorming ideas, prototyping, and testing solutions. Widely applied in fields such as engineering and business, design thinking fosters innovation by blending empathy, experimentation, and practical implementation [71, 54]. In the context of school education, Li et al. argue that design thinking should be regarded as a model of thinking to enhance students’ learning [74, p.97]. They cite several studies showing how students can learn through design based activities and also develop design thinking skills [74, p.98]. A valuable resource for educators is the Educational Designer [75], an international open-access e-journal focused on research-based practices in the design, development, and evaluation of educational materials in mathematics, science, engineering, and technology[74, p.101].
@@
The College Board Standards for College Success [55, p.5] outlines five core science practices that shape scientific inquiry. These include: (i) formulating empirically testable questions and predictions, (ii) collecting relevant data to address these questions, (iii) identifying patterns in data through analysis, (iv) constructing scientific explanations using evidence and knowledge, and (v) applying mathematical reasoning to interpret data and solve problems. These practices emphasize critical thinking and methodological rigor in science education.
@@
The Framework for K-12 Science Education [79],  [43, p.xv] identifies three essential dimensions for providing a high-quality science education: Science and Engineering Practice, which engage students in scientific inquiry and problem-solving; Crosscutting Concepts, which include overarching themes such as patterns and systems that apply across disciplines; and Disciplinary Core Ideas, which focus on the key ideas in physical, life, earth, space, and other sciences. Together, these dimensions help students not only understand scientific content but also grasp how it is acquired and the connections that exist between various scientific disciplines.
@@
The Next Generation Science Standards 2013 document also emphasizes the importance of “instructional flexibility” [43, p.xiv], empowering educators to tailor their approaches based on local contexts to enhance student learning. This flexibility is crucial for integrating disciplines such as mathematics and computation, which naturally intersect with scientific inquiry.
@@
The AAPT Recommendations for the Undergraduate Physics Laboratory Curriculum [56, p.2,3] report, published by the American Association of Physics Teachers (AAPT), emphasizes promoting physics-based thinking through several key focus areas. (i) Constructing knowledge: Students are encouraged to independently collect, analyze, and interpret real data, building confidence in their ability to generate scientific insights. (ii) Modeling: Developing abstract representations of physical systems is vital, enabling students to predict outcomes, interpret results, and recognize model limitations. (iii) Designing Experiments: Students need to be guided to ask scientific questions, engineer solutions, and troubleshoot, gaining valuable hands-on experience throughout the process. (iv) Developing Technical and Practical Laboratory Skills: Proficiency with laboratory equipment and understanding the appropriate tools for different measurements are key priorities. (v) Analyzing and Visualizing Data: Students are expected to apply statistical methods, quantify uncertainties, and effectively represent findings through data visualization. (vi) Communicating Physics: Students learn to present scientific results in authentic formats, such as reports and presentations, while developing teamwork and ethical communication skills. Together, these focus areas aim to cultivate a comprehensive set of skills essential for scientific inquiry.
@@
In the ED problem shown in Figure 2, students’ science-based thinking (or, science thinking) may include constructing knowledge through the collection and analysis of real data related to delivery mechanisms such as ziplines, drones, or catapults. Alternatively, they may apply their physics knowledge to propose solutions. Each approach may involve distinct scientific concepts or ideas: ziplines might require an understanding of tension and gravitational forces, while drones introduce concepts of aerodynamics and propulsion. Students might also consider applying the momentum principle [34] to evaluate how the payload’s momentum changes throughout its journey, ensuring stability and precision. When working with catapults, they may analyze projectile motion, accounting for factors such as launch angle, initial velocity, and air drag to achieve accurate delivery. The energy principle [34] can also come into play, as students explore the conversion of potential energy into kinetic energy, optimizing their designs for maximum range and effectiveness. Rather than unfolding in isolated silos, design thinking and science thinking appear naturally interwoven throughout these activities. This interplay, revealed in our analysis of student data, motivated us to advocate for the usage of ‘design-science connection’ as opposed to ‘design-science gap’ in our recent study [21].
@@
The Committee on the Undergraduate Program in Mathematics in its 2015 report, emphasizes the need to foster a “mathematical habits of mind” [58, p.10,11] by developing students’ cognitive abilities and content knowledge. For cognitive goals, the committee recommends: (i) developing students’ thinking and communication skills to reason logically and express mathematical ideas clearly, (ii) linking mathematical theory with interdisciplinary real-world applications, (iii) using technological tools effectively for problem-solving and exploration, and (iv) fostering mathematical independence through open-ended inquiry. As for the content, the committee recommends undergraduate programs should cover core areas such as calculus, linear algebra, data analysis, computing, and mathematical modeling. A similar perspective is articulated in the 2013 report by the Committee on the Mathematical Sciences in 2025 [80], which emphasizes the need to adapt undergraduate education to meet future demands in research, industry, and interdisciplinary applications.
@@
In our context, mathematics serves as an indispensable tool for solving engineering design-based problems. This perspective is further reinforced by Tiles’ assertion that mathematics “is not just a language which can be translated into and out of, rather it plays an essential role in the articulation of scientific concepts” [59, p.3].
@@
The National Research Council’s How People Learn: Brain, Mind, Experience, and School: Expanded Edition (2000) [62] serves as a foundational document that underscores the significance of metacognition in student learning. The report argues that adopting a metacognitive approach to instruction “can help students learn to take control of their own learning by defining learning goals and monitoring their progress in achieving them” [62, p.18]. This perspective is built on the understanding that metacognitive thinking encourages students to engage in an “internal conversation” [62, p.18] that fosters self-regulation and self-directed learning. Furthermore, it posits that metacognition can empower students to “teach themselves” [62, p.50]. Consequently, the report advocates for the integration of metacognitive skills into various subject areas within the curriculum.
@@
In the context of this report, the lead author ‘reflected’ on the definition of metacognition. Initially, he perceived it simply as “reflecting, or thinking about thinking”. However, he realized this may be a naive view considering Crawford and Capps’ assertion that “Reflection and metacognition are both thinking processes. Yet, reflection and metacognition differ in some ways” [63, p.16]. Our aim in this article is not to engage in a debate regarding the precise definition of metacognition. Rather, it is perhaps prudent to align with Livingston’s assertion that “In actuality, defining metacognition is not that simple” [82, p.2]. Given the complexities surrounding the term, we find it beneficial to draw on the fine example provided by Kohen and Kramarski, who note that “presenting data in a graph is a cognitive function, whereas reflecting on the answer and realizing that the graph fits the givens are part of the metacognitive process” [63, p.280]. For the records, we consider metacognition as a “deliberate, planful, and goal-oriented mental process, using higher level thinking skills applied to one’s thoughts and experiences” [63, p.17].
@@
A compelling reason for incorporating metacognitive reflection as a critical way of thinking within our framework is best articulated in Hacker and Dunlosky’s thoughtful essay, Not All Metacognition Is Created Equal. They emphasize that students “should not be instructed to merely use ‘metacognition’ but instead should be directed toward specific, higher-level analyses in which they explain how they are solving the problem” [64, p.79]. This insistence on directing students toward higher-level analysis highlights the need for metacognitive reflection to be not only present, but also be purposeful in guiding students toward more profound understanding and effective problem-solving strategies.
@@
A similar perspective is echoed in the NRC’s 2012 report, which highlights how metacognitive approaches in STEM education are embedded in problem-based learning and reflective activities designed to promote critical thinking. While instructors may assume students possess these skills or view them as too advanced for introductory courses, research shows that explicitly teaching metacognitive skills alongside science content enhances student responsibility, problem-solving abilities, and study habits [44, p.154].
@@
Another recent study by Ulu and Yerdelen-Damar [86, p.20, 21] employs a metacognition framework inspired by Brown, Schraw, and Dennison [87, 88]. Metacognition refers to understanding and regulating one’s own thinking and is divided into two components: (i) Knowledge of cognition—declarative information about one’s cognitive processes—and (ii) Regulation of cognition—the ability to plan, monitor, control, and evaluate one’s thinking. Knowledge of cognition consists of three levels: declarative knowledge (facts and strategies), procedural knowledge (how to apply strategies), and conditional knowledge (when and why to apply strategies). Regulation of cognition involves five key skills: planning, managing information, monitoring comprehension, debugging, and evaluating. Although we did not adopt this framework in our earlier studies, it holds promise for future integration into our WoT4EDP framework. Its structured approach to analyzing metacognitive thinking aligns with the reflective practices we aim to foster in design-based activities.
@@
Integrating computational practices within STEM classrooms equips learners with an authentic understanding of scientific processes and better prepares them for careers in an increasingly computational world. As computational thinking (CT) becomes essential across STEM disciplines, the demand for computation-integrated physics courses continues to grow, underscoring the notion that computation is synonymous with engaging in STEM. The urgency and significance of incorporating computational thinking (CT) based activities into physics education is powerfully highlighted by Hamerski et al. [89, p.23], who raise a “call to action” in their case study of high school students involved in a computation-integrated physics course. They urge researchers to recognize and respond to the pressing need for integrating computation into educational practices.
@@
While educators may choose various computational thinking (CT) frameworks [67, p.151, 152] that best suit their contexts, we have opted for Shute et al.’s framework due to its relevance to our instructional goals. Our students work extensively with Python in the Jupyter Notebook [35] environment via Google Colab [36], which makes Shute et al.’s definition of CT particularly relevant, simple, and practical: “CT as the conceptual foundation required to solve problems effectively and efficiently (i.e., algorithmically, with or without the assistance of computers), with solutions that are reusable in different contexts” [67, p.153]. This CT framework comprises six key facets: (i) Decomposition—breaking a complex system or problem into manageable, functional sub-problems; (ii) Abstraction—extracting key elements, including data collection, pattern recognition, and modeling; (iii) Algorithms—developing logical, step-by-step instructions for solving problems, often represented through flowcharts and implemented in code (e.g., Python); (iv) Debugging—identifying and fixing errors in code or algorithms; (v) Iteration—refining the algorithm or code through repeated improvements; and (vi) Generalization—applying algorithms, code, or CT skills across different domains, enabling efficient problem-solving [67, p.153].
Importantly, Shute et al. emphasize that this framing of CT underpins and is evidenced through a variety of subject areas, “including Mathematics, Science, and even English Language Arts”. This perspective reinforces how various ways of thinking—mathematical, scientific, and computational—are closely linked, facilitating a holistic approach to problem-solving in our educational practices [67, p.146, 151].
@@ 3.5 Pedagogical Goals
Problem Solving is a most natural goal for any science educator, yet it is almost taken for granted. The AAPT Recommendations for the Undergraduate Physics Laboratory Curriculum emphasize that physics is an approach to problem-solving that relies on direct observation and hands-on experimentation. To succeed in this field, one must integrate a wide range of knowledge and skills, including mathematical, computational, experimental, and practical abilities, while also cultivating a “thinking like a physicist” mindset [56, p.iii]. A strong grasp of scientific principles is important for students to effectively and efficiently solve problems. As the Next Generation Science Standards (NGSS) 2013 report rightly notes, “higher-level reasoning and problem-solving practices require a reasonable depth of familiarity with the content of a given scientific topic if students are to engage in them meaningfully” [43, p.15].
Regarding Learning Innovation, we are largely guided by English’s work, and we will return to this in more detail in Section (iv). While innovation in this context refers to students’ learning, we are tempted to extend it as a goal for educators too. We would approach the term ‘innovation’ rather cautiously, and the SPIN-UP Report aptly captures our tempered view: true ‘Eureka’ moments are rare in science, and even rarer in education. Progress in teaching and learning is incremental and ongoing, with innovations needing to be adapted to suit local needs [100, p.49, 67].

### v4

@@ 3.4 The Five WoT4EDP Elements
DBT Economic aspects, safety aspects, criteria and constraints, identify stakeholders, contraption dimensions, contraption mechanism, design limitations, feasibility study. [20, 21] Design thinking helps students develop problem-solving skills, creativity, empathy, collaboration, iterative thinking, and multiple perspectives. [57] Incorporate real-world, open-ended problems to encourage creative and user-centered solutions. Use iterative prototyping and reflective feedback sessions to help students refine ideas and increase empathy. Collaborate across disciplines to stimulate innovation. A focus on failure as a learning opportunity strengthens resilience. The ‘design thinking competency model’ [58, p.342] may be used as a guide.
@@
Based on our readings [71, p.4], we gather that the term design thinking likely originated with Rowe, who used it as the title of his 1987 book [72], in which he aimed to provide a structured account of problem-solving processes in design. Since then, however, the meaning of the term design thinking has undergone several transformations and has drawn varied interpretations.
Haasi and Laakso [71, p.5] describe design thinking as comprising three key elements: (i) practices, including human-centered approaches, thinking by doing, visualizing, integrating multiple approaches, and collaboration; (ii) cognitive approaches, such as abductive reasoning, reflective reasoning, a holistic perspective, and integrative thinking; and (iii) a mindset characterized by exploration, tolerance for ambiguity, optimism, and a future-oriented outlook.
@@
While the debate over the usage of design thinking is thought-provoking, we find merit in the use of the term in our contexts. For operational purposes, one may view design thinking as a problem-solving approach that emphasizes creativity, user-centered solutions, and iterative processes. It involves understanding users’ needs, brainstorming ideas, prototyping, and testing solutions. Widely applied in fields such as engineering and business, design thinking fosters innovation by blending empathy, experimentation, and practical implementation [73, 58]. In the context of school education, Li et al. argue that design thinking should be regarded as a model of thinking to enhance students’ learning [76, p.97]. They cite several studies showing how students can learn through design based activities and also develop design thinking skills [76, p.98]. A valuable resource for educators is the Educational Designer [77], an international open-access e-journal focused on research-based practices in the design, development, and evaluation of educational materials in mathematics, science, engineering, and technology[76, p.101].
@@
The College Board Standards for College Success [59, p.5] outlines five core science practices that shape scientific inquiry. These include: (i) formulating empirically testable questions and predictions, (ii) collecting relevant data to address these questions, (iii) identifying patterns in data through analysis, (iv) constructing scientific explanations using evidence and knowledge, and (v) applying mathematical reasoning to interpret data and solve problems. These practices emphasize critical thinking and methodological rigor in science education.
@@
The Framework for K-12 Science Education [24],  [23, p.xv] identifies three essential dimensions for providing a high-quality science education: Science and Engineering Practice, which engage students in scientific inquiry and problem-solving; Crosscutting Concepts, which include overarching themes such as patterns and systems that apply across disciplines; and Disciplinary Core Ideas, which focus on the key ideas in physical, life, earth, space, and other sciences. Together, these dimensions help students not only understand scientific content but also grasp how it is acquired and the connections that exist between various scientific disciplines.
@@
The Next Generation Science Standards 2013 document also emphasizes the importance of “instructional flexibility” [23, p.xiv], empowering educators to tailor their approaches based on local contexts to enhance student learning. This flexibility is crucial for integrating disciplines such as mathematics and computation, which naturally intersect with scientific inquiry.
@@
The AAPT Recommendations for the Undergraduate Physics Laboratory Curriculum [22, p.2,3] report, published by the American Association of Physics Teachers (AAPT), emphasizes promoting physics-based thinking through several key focus areas. (i) Constructing knowledge: Students are encouraged to independently collect, analyze, and interpret real data, building confidence in their ability to generate scientific insights. (ii) Modeling: Developing abstract representations of physical systems is vital, enabling students to predict outcomes, interpret results, and recognize model limitations. (iii) Designing Experiments: Students need to be guided to ask scientific questions, engineer solutions, and troubleshoot, gaining valuable hands-on experience throughout the process. (iv) Developing Technical and Practical Laboratory Skills: Proficiency with laboratory equipment and understanding the appropriate tools for different measurements are key priorities. (v) Analyzing and Visualizing Data: Students are expected to apply statistical methods, quantify uncertainties, and effectively represent findings through data visualization. (vi) Communicating Physics: Students learn to present scientific results in authentic formats, such as reports and presentations, while developing teamwork and ethical communication skills. Together, these focus areas aim to cultivate a comprehensive set of skills essential for scientific inquiry.
@@
In the ED problem shown in Figure 2, students’ science-based thinking (or, science thinking) may include constructing knowledge through the collection and analysis of real data related to delivery mechanisms such as ziplines, drones, or catapults. Alternatively, they may apply their physics knowledge to propose solutions. Each approach may involve distinct scientific concepts or ideas: ziplines might require an understanding of tension and gravitational forces, while drones introduce concepts of aerodynamics and propulsion. Students might also consider applying the momentum principle [40] to evaluate how the payload’s momentum changes throughout its journey, ensuring stability and precision. When working with catapults, they may analyze projectile motion, accounting for factors such as launch angle, initial velocity, and air drag to achieve accurate delivery. The energy principle [40] can also come into play, as students explore the conversion of potential energy into kinetic energy, optimizing their designs for maximum range and effectiveness. Rather than unfolding in isolated silos, design thinking and science thinking appear naturally interwoven throughout these activities. This interplay, revealed in our analysis of student data, motivated us to advocate for the usage of design-science connection as opposed to design-science gap in our recent study [21].
@@
The Committee on the Undergraduate Program in Mathematics in its 2015 report, emphasizes the need to foster a “mathematical habits of mind” [26, p.10,11] by developing students’ cognitive abilities and content knowledge. For cognitive goals, the committee recommends: (i) developing students’ thinking and communication skills to reason logically and express mathematical ideas clearly, (ii) linking mathematical theory with interdisciplinary real-world applications, (iii) using technological tools effectively for problem-solving and exploration, and (iv) fostering mathematical independence through open-ended inquiry. As for the content, the committee recommends undergraduate programs should cover core areas such as calculus, linear algebra, data analysis, computing, and mathematical modeling. A similar perspective is articulated in the 2013 report by the Committee on the Mathematical Sciences in 2025 [81], which emphasizes the need to adapt undergraduate education to meet future demands in research, industry, and interdisciplinary applications.
@@
In our context, mathematics serves as an indispensable tool for solving engineering design-based problems. This perspective is further reinforced by Tiles’ assertion that mathematics “is not just a language which can be translated into and out of, rather it plays an essential role in the articulation of scientific concepts” [61, p.3].
@@
The National Research Council’s How People Learn: Brain, Mind, Experience, and School: Expanded Edition (2000) [64] serves as a foundational document that underscores the significance of metacognition in student learning. The report argues that adopting a metacognitive approach to instruction “can help students learn to take control of their own learning by defining learning goals and monitoring their progress in achieving them” [64, p.18]. This perspective is built on the understanding that metacognitive thinking encourages students to engage in an “internal conversation” [64, p.18] that fosters self-regulation and self-directed learning. Furthermore, it posits that metacognition can empower students to “teach themselves” [64, p.50]. Consequently, the report advocates for the integration of metacognitive skills into various subject areas within the curriculum.
@@
In the context of this report, the lead author ‘reflected’ on the definition of metacognition. Initially, he perceived it simply as “reflecting, or thinking about thinking”. However, he realized this may be a naive view considering Crawford and Capps’ assertion that “Reflection and metacognition are both thinking processes. Yet, reflection and metacognition differ in some ways” [65, p.16]. Our aim in this article is not to engage in a debate regarding the precise definition of metacognition. Rather, it is perhaps prudent to align with Livingston’s assertion that “In actuality, defining metacognition is not that simple” [83, p.2]. Given the complexities surrounding the term, we find it beneficial to draw on the fine example provided by Kohen and Kramarski, who note that “presenting data in a graph is a cognitive function, whereas reflecting on the answer and realizing that the graph fits the givens are part of the metacognitive process” [65, p.280]. For the records, we consider metacognition as a “deliberate, planful, and goal-oriented mental process, using higher level thinking skills applied to one’s thoughts and experiences” [65, p.17].
@@
A compelling reason for incorporating metacognitive reflection as a critical way of thinking within our framework is best articulated in Hacker and Dunlosky’s thoughtful essay, Not All Metacognition Is Created Equal. They emphasize that students “should not be instructed to merely use ‘metacognition’ but instead should be directed toward specific, higher-level analyses in which they explain how they are solving the problem” [66, p.79]. This insistence on directing students toward higher-level analysis highlights the need for metacognitive reflection to be not only present, but also be purposeful in guiding students toward more profound understanding and effective problem-solving strategies.
@@
A similar perspective is echoed in the NRC’s 2012 report, which highlights how metacognitive approaches in STEM education are embedded in problem-based learning and reflective activities designed to promote critical thinking. While instructors may assume students possess these skills or view them as too advanced for introductory courses, research shows that explicitly teaching metacognitive skills alongside science content enhances student responsibility, problem-solving abilities, and study habits [25, p.154].
@@
Another recent study by Ulu and Yerdelen-Damar [87, p.20, 21] employs a metacognition framework inspired by Brown, Schraw, and Dennison [88, 89]. Metacognition refers to understanding and regulating one’s own thinking and is divided into two components: (i) Knowledge of cognition—declarative information about one’s cognitive processes—and (ii) Regulation of cognition—the ability to plan, monitor, control, and evaluate one’s thinking. Knowledge of cognition consists of three levels: declarative knowledge (facts and strategies), procedural knowledge (how to apply strategies), and conditional knowledge (when and why to apply strategies). Regulation of cognition involves five key skills: planning, managing information, monitoring comprehension, debugging, and evaluating. Although we did not adopt this framework in our earlier studies, it holds promise for future integration into our WoT4EDP framework. Its structured approach to analyzing metacognitive thinking aligns with the reflective practices we aim to foster in design-based activities.
@@
Integrating computational practices within STEM classrooms equips learners with an authentic understanding of scientific processes and better prepares them for careers in an increasingly computational world. As computational thinking (CT) becomes essential across STEM disciplines, the demand for computation-integrated physics courses continues to grow, underscoring the notion that computation is synonymous with engaging in STEM. The urgency and significance of incorporating computational thinking (CT) based activities into physics education is powerfully highlighted by Hamerski et al. [90, p.23], who raise a “call to action” in their case study of high school students involved in a computation-integrated physics course. They urge researchers to recognize and respond to the pressing need for integrating computation into educational practices.
@@
While educators may choose various computational thinking (CT) frameworks [69, p.151, 152] that best suit their contexts, we have opted for Shute et al.’s framework due to its relevance to our instructional goals. Our students work extensively with Python in the Jupyter Notebook [41] environment via Google Colab [42], which makes Shute et al.’s definition of CT particularly relevant, simple, and practical: “CT as the conceptual foundation required to solve problems effectively and efficiently (i.e., algorithmically, with or without the assistance of computers), with solutions that are reusable in different contexts” [69, p.153]. This CT framework comprises six key facets: (i) Decomposition—breaking a complex system or problem into manageable, functional sub-problems; (ii) Abstraction—extracting key elements, including data collection, pattern recognition, and modeling; (iii) Algorithms—developing logical, step-by-step instructions for solving problems, often represented through flowcharts and implemented in code (e.g., Python); (iv) Debugging—identifying and fixing errors in code or algorithms; (v) Iteration—refining the algorithm or code through repeated improvements; and (vi) Generalization—applying algorithms, code, or CT skills across different domains, enabling efficient problem-solving [69, p.153].
Importantly, Shute et al. emphasize that this framing of CT underpins and is evidenced through a variety of subject areas, “including Mathematics, Science, and even English Language Arts”. This perspective reinforces how various ways of thinking—mathematical, scientific, and computational—are closely linked, facilitating a holistic approach to problem-solving in our educational practices [69, p.146, 151].
@@ 3.5 Pedagogical Goals
Problem Solving is a most natural goal for any science educator, yet it is almost taken for granted. The AAPT Recommendations for the Undergraduate Physics Laboratory Curriculum emphasize that physics is an approach to problem-solving that relies on direct observation and hands-on experimentation. To succeed in this field, one must integrate a wide range of knowledge and skills, including mathematical, computational, experimental, and practical abilities, while also cultivating a “thinking like a physicist” mindset [22, p.iii]. A strong grasp of scientific principles is important for students to effectively and efficiently solve problems. As the Next Generation Science Standards (NGSS) 2013 report rightly notes, “higher-level reasoning and problem-solving practices require a reasonable depth of familiarity with the content of a given scientific topic if students are to engage in them meaningfully” [23, p.15].
Regarding Learning Innovation, we are largely guided by English’s work, and we will return to this in more detail in Section (iv). While innovation in this context refers to students’ learning, we are tempted to extend it as a goal for educators too. We would approach the term innovation rather cautiously, and the SPIN-UP Report aptly captures our tempered view: true Eureka moments are rare in science, and even rarer in education. Progress in teaching and learning is incremental and ongoing, with innovations needing to be adapted to suit local needs [100, p.49, 67].